# GraspCoT: Integrating Physical Property Reasoning for 6-DoF Grasping under Flexible Language Instructions

### v1

@@ Abstract
Flexible instruction-guided 6-DoF grasping is a significant yet challenging task for real-world robotic systems. Existing methods utilize the contextual understanding capabilities of the large language models (LLMs) to establish mappings between expressions and targets, allowing robots to comprehend users’ intentions in the instructions. However, the LLM’s knowledge about objects’ physical properties remains underexplored despite its tight relevance to grasping. In this work, we propose GraspCoT, a 6-DoF grasp detection framework that integrates a Chain-of-Thought (CoT) reasoning mechanism oriented to physical properties, guided by auxiliary question-answering (QA) tasks. Particularly, we design a set of QA templates to enable hierarchical reasoning that includes three stages: target parsing, physical property analysis, and grasp action selection. Moreover, GraspCoT presents a unified multimodal LLM architecture, which encodes multi-view observations of 3D scenes into 3D-aware visual tokens, and then jointly embeds these visual tokens with CoT-derived textual tokens within LLMs to generate grasp pose predictions. Furthermore, we present IntentGrasp, a large-scale benchmark that fills the gap in public datasets for multi-object grasp detection under diverse and indirect verbal commands. Extensive experiments on IntentGrasp demonstrate the superiority of our method, with additional validation in real-world robotic applications confirming its practicality. Codes and data will be released.
@@ 2 Related Work
Robotic Grasping. The emergence of grasping datasets [8, 7, 25, 5] lay the groundwork for data-driven grasping, while later works extend these foundations to diverse and task-aware scenarios. 6-DOF GraspNet [26] employs variational autoencoders to sample diverse poses from partial observations, and AnyGrasp [9] introduces temporal consistency for dynamic environments. Recognizing the need for hardware generalization, UniGrasp [35] pioneers gripper-agnostic frameworks through kinematic latent encoding, while optimization-centric paradigms like SE(3)-DiffusionFields [40] reformulate grasping as gradient-based pose-trajectory co-optimization. Recent efforts balance these directions: Wang et al. [41] accelerates inference via geometric heuristics, and Wang et al. [42] enhances physical interaction through auxiliary task learning. Collectively, these works demonstrate a paradigm shift from isolated geometric analysis to hybrid systems that embed functional adaptability and computational efficiency within unified frameworks.
@@
Language-guided Grasping. Recent advances in vision-language models have revolutionized language-guided robotic grasping by enabling semantic alignment between textual instructions and visual perception. Pioneering works [29, 30, 36, 28, 19, 49, 23] leverage a text backbone [33, 6] to embed textual features, where some explore diffusion models for 3D manipulation–3DAPNet [29] extends grasp pose prediction paradigm to open-vocabulary affordance-pose joint learning, while LGrasp6D [30] introduces the Grasp-Anything-6D dataset and incorporates the negative prompt guidance for 6-DoF grasp detection. OVGNet [19] proposes a unified visual-linguistic framework combining CLIP-style cross-modal alignment with grasp pose prediction, enabling generalization to unseen categories without category-specific retraining. The emergence of (M)LLMs [38, 22, 17] further expands the visual and textual interpretation capabilities in manipulation tasks [37, 18, 39]. ThinkGrasp [31] introduces a language-guided, stepwise clutter removal strategy powered by GPT-4o’s contextual inferring, enabling robotic grasping in occluded environments through goal-oriented obstacle displacement and few-step target recovery. RT-Grasp [47] proposes a framework that unlocks LLMs’ capacity for numerical prediction in robotics by embedding structured inferring phases into training, while Jin et al. [13] leverage an end-to-end multimodal LLM framework to bridge implicit human intentions to grasp poses, enabling object/part-level grasping under indirect instructions. Recent efforts combine visual grounding with attribute prediction of required objects using multimodal foundation models. ShapeGrasp [20] pioneers zero-shot task-oriented grasping through geometric decomposition of objects into convex shapes represented as attribute-rich graphs, combined with LLM-driven semantic reasoning to infer part-task utility. DeliGrasp [46] proposes LLM-driven object properties prediction for adaptive grasping, converting semantic object descriptions into grasp policies via inferred mass/friction/spring parameters. Instead of requiring LLMs to directly predict precise physical parameters–a task challenging current LLMs’ capabilities–our method encodes physical properties into structured descriptors, explicitly guiding grasp pose prediction while avoiding error-prone low-level physics estimation.
@@ 3.3 Physical-Property-Oriented CoT Design
    [<$\texttt{obj}_{1}$>, <$\texttt{obj}_{2}$>, …]. Total <#objs>.
@@
    The material properties of <$\texttt{obj}_{\texttt{k}}$> are <$\delta_{\texttt{phy}}$> (hardness), <$\delta_{\texttt{phy}}$> (strength), and <$\delta_{\texttt{phy}}$> (elasticity).
@@
    The appropriate verb to grasp the <$\texttt{obj}_{\texttt{k}}$> is <$\delta_{\texttt{act}}$>.
@@ 3.5 IntentGrasp Benchmark Construction
Table 1: Results on IntentGrasp benchmark under flexible language instructions. “CR@$\theta$” denotes the success rate when predicted poses are within a threshold $\theta$ of ground truth poses.
@@ 4.1 Main Results on IntentGrasp Benchmark
Setup. To evaluate the effectiveness of our grasp pose prediction method, we employ three metrics: coverage rate (CR) [26], earth mover’s distance (EMD) [40], and collision-free rate (CFR) [52] with appropriate adaptations. CR quantifies the proportion of ground-truth grasp poses that are sufficiently covered by predictions. Specifically, it measures the percentage of ground-truth poses where at least one predicted pose lies within a distance threshold $\theta$ (Euclidean distance in SE(3) space). We evaluate CR under three increasingly stringent thresholds: $\theta$ = 0.4, 0.3, and 0.2. EMD measures how holistically the predicted poses match the spatial distribution of ground-truth annotations in SE(3) space, where lower values indicate better distributional consistency. CFR evaluates the physical feasibility of predicted grasp poses by measuring their ability to avoid collisions with the target object and environment. However, a critical limitation arises when directly using the CFR metric: erroneous grasp poses far from the target object may artificially achieve higher scores because they avoid contact with environmental point clouds entirely. To address this, we propose an EMD-Weighted CFR (EW-CFR) metric that jointly evaluates geometric consistency and physical feasibility where higher values indicate safer grasps. The IntentGrasp dataset is partitioned following conventional practice, with 80% of samples dedicated to training and the remaining 20% preserved as the evaluation set.
@@
Grasp Label Pruning. Wu [45] et al. demonstrate that excessively dense supervision incurs label ambiguity, particularly caused by the diversity of grasp orientations, as a critical bottleneck hindering model convergence. Our analysis reveals that certain objects in the Grasp-Anything-6D dataset contain over 700 annotated grasp poses. This overabundance of positive samples not only exacerbates the mentioned issues but also severely challenges Hungarian matching during training. Consequently, we implemented a clustering pruning strategy in IntentGrasp to reduce similar grasp labels, especially redundant rotation-similar annotations, per object to 100 instances. The detailed pruning method is provided in the supplementary material.
Training. Our models are trained using the Adam optimizer with a batch size of 8 per GPU over 4 epochs, initialized with LLaVA-3D [54] pre-trained weights from the HuggingFace repository. The learning rate followed a cosine annealing schedule, starting at 1e-4. All experiments are executed on 8$\times$24GB NVIDIA RTX3090 GPUs.
@@
Quantitative Results. In Tab. 1, our method presents superior performance across all metrics on the IntentGrasp benchmark under flexible language instructions. Notably, GraspCoT attains a CR of 0.5587 at the strictest threshold ($\theta$=0.2), outperforming 3DAPNet and LGrasp6D by a large margin of 0.2683 and 0.2238, respectively. This substantial improvement highlights our method’s ability to generate geometrically precise grasp poses even under high-precision requirements. The progressive performance gap across $\theta$=0.4, 0.3, and 0.2 further demonstrates that our approach maintains effectiveness as task difficulty escalates. The EMD of GraspCoT reflects a 14.1% relative reduction compared to LGrasp6D, indicating superior alignment between the predicted and ground-truth pose distributions. For our proposed EW-CFR metric, GraspCoT achieves a score of 0.4229, surpassing 3DAPNet and LGrasp6D by 20.1% and 9.9%, respectively. The w/o CoT version of GraspCoT confirms the critical role of our CoT reasoning pipeline: removing CoT results in a 25.4% decline in CR ($\theta$=0.2) and a 9.1% increase in EMD, respectively.
@@ 4.2 Ablation Studies
4.3 Real-World Experiment
(a) Real-world experimental setup.
(b) Example of the grasping task.
Figure 6: An application example on a Kinova Gen3 robotic arm.
Setup. We employ a Kinova Gen3 robotic arm equipped with an Intel RealSense D435i depth camera for real-world deployment. The camera detects 6-DoF grasp poses, which are mapped to the robot’s end-effector poses using transformation matrices obtained through hand-eye calibration. The Kinova arm automatically plans and executes the trajectory based on the received poses, ensuring a precise and efficient grasp execution. Our platform consists of a computer equipped with an NVIDIA RTX 3090, running the Robot Operating System (ROS). The ROS is utilized to drive the camera for real-time data acquisition and to execute the Kinova robotic arm driver program, which communicates with the Kinova robot via a network interface card (NIC). The inference program, based on ROS, receives RGB-D data from the camera, performs inference, and outputs the detected 6-DoF grasp poses, which are then converted and sent to the Kinova robot for execution. We categorize the test scenarios into explicit instructions and flexible instructions, further dividing the latter into single-target grasping and multi-target sequential grasping. For each setting, we conduct 35 repeated grasping experiments, with the success rate as the evaluation metric. Partial successes (e.g., grasping one of two required targets) are counted as individual successes, with the rate defined as successful grasps / total targets.
Results. As shown in Tab. 5, when handling explicit grasping instructions, the CoT-enhanced GraspCoT achieved a 54.2% success rate, outperforming its non-CoT counterpart by 2.8%. For flexible instructions, the CoT version maintained stable performance with a 5.6% advantage over the non-CoT variant, further validating its advantages in optimizing the grasp pose targeted at the object itself. Notably in multi-target grasping scenarios where avoiding omissions of targets with weaker implicit relevance remains challenging, the CoT implementation improved success rates from 44.0% to 46.7%, which is a 2.7% enhancement, verifying its effectiveness in handling complex requirements.
@@ 5 Conclusion and Discussion
This work presents a novel 6-DoF grasp detection framework that bridges the gap between flexible instructions and robotic grasping. By integrating a physical-property-oriented CoT mechanism with QA-guided hierarchical reasoning, our method aligns target parsing, physical property analysis, and grasp action selection with grasp detection. The proposed multimodal LLM architecture, augmented by 3D-aware visual-textual token fusion, achieves state-of-the-art performance on a newly established benchmark IntentGrasp for diverse verbal-command grasping. Real-world validations further underscore its practical applicability. While our method demonstrates robust performance under flexible instructions, some limitations persist. For instance, grasp predictions for slender objects (e.g., chopsticks) occasionally deviate from mass centers despite leveraging shape descriptors, suggesting the need to model latent physical properties beyond geometric cues. Additionally, future directions include extending the framework to sequential video inputs for dynamic scenes, such as grasping moving targets, which would require video-based grasping benchmarks to advance real-world adaptability.

### v2

@@ Abstract
Flexible instruction-guided 6-DoF grasping is a significant yet challenging task for real-world robotic systems. Existing methods utilize the contextual understanding capabilities of the large language models (LLMs) to establish mappings between expressions and targets, allowing robots to comprehend users’ intentions in the instructions. However, the LLM’s knowledge about objects’ physical properties remains underexplored despite its tight relevance to grasping. In this work, we propose GraspCoT, a 6-DoF grasp detection framework that integrates a Chain-of-Thought (CoT) reasoning mechanism oriented to physical properties, guided by auxiliary question-answering (QA) tasks. Particularly, we design a set of QA templates to enable hierarchical reasoning that includes three stages: target parsing, physical property analysis, and grasp action selection. Moreover, GraspCoT presents a unified multimodal LLM architecture, which encodes multi-view observations of 3D scenes into 3D-aware visual tokens, and then jointly embeds these visual tokens with CoT-derived textual tokens within LLMs to generate grasp pose predictions. Furthermore, we present IntentGrasp, a large-scale benchmark that fills the gap in public datasets for multi-object grasp detection under diverse and indirect verbal commands. Extensive experiments on IntentGrasp demonstrate the superiority of our method, with additional validation in real-world robotic applications confirming its practicality. The code is available at https://github.com/cxmomo/GraspCoT.
@@ 2 Related Work
Robotic Grasping. The emergence of grasping datasets [8, 7, 25, 5] lay the groundwork for data-driven grasping, while later works extend these foundations to diverse and task-aware scenarios. 6-DOF GraspNet [26] employs variational autoencoders to sample diverse poses from partial observations, and AnyGrasp [9] introduces temporal consistency for dynamic environments. Recognizing the need for hardware generalization, UniGrasp [36] pioneers gripper-agnostic frameworks through kinematic latent encoding, while optimization-centric paradigms like SE(3)-DiffusionFields [41] reformulate grasping as gradient-based pose-trajectory co-optimization. Recent efforts balance these directions: Wang et al. [42] accelerates inference via geometric heuristics, and Wang et al. [43] enhances physical interaction through auxiliary task learning. Collectively, these works demonstrate a shift from isolated geometric analysis to hybrid systems that embed functional adaptability and computational efficiency within unified frameworks.
@@
Language-guided Grasping. Recent advances in vision-language models have revolutionized language-guided robotic grasping by enabling semantic alignment between textual instructions and visual perception. Pioneering works [29, 30, 37, 28, 19, 51, 23, 46] leverage a text backbone [34, 6] to embed textual features, where some explore diffusion models for 3D manipulation–3DAPNet [29] extends grasp pose prediction paradigm to open-vocabulary affordance-pose joint learning, while LGrasp6D [30] introduces the Grasp-Anything-6D dataset and incorporates the negative prompt guidance for 6-DoF grasp detection. OVGNet [19] proposes a unified visual-linguistic framework combining CLIP-style cross-modal alignment with grasp pose prediction, enabling generalization to unseen categories without category-specific retraining. The emergence of (M)LLMs [39, 22, 17] further expands the visual and textual interpretation capabilities in manipulation tasks [38, 18, 40, 31]. ThinkGrasp [32] introduces a language-guided, stepwise clutter removal strategy powered by GPT-4o’s contextual inferring, enabling robotic grasping in occluded environments through goal-oriented obstacle displacement and few-step target recovery. RT-Grasp [49] proposes a framework that unlocks LLMs’ capacity for numerical prediction in robotics by embedding structured inferring phases into training, while Jin et al. [13] leverage an end-to-end multimodal LLM framework to bridge implicit human intentions to grasp poses, enabling object/part-level grasping under indirect instructions. Recent efforts combine visual grounding with attribute prediction of required objects using multimodal foundation models. ShapeGrasp [20] pioneers zero-shot task-oriented grasping through geometric decomposition of objects into convex shapes represented as attribute-rich graphs, combined with LLM-driven semantic reasoning to infer part-task utility. DeliGrasp [48] proposes LLM-driven object properties prediction for adaptive grasping, converting semantic object descriptions into grasp policies via inferred mass/friction/spring parameters. Instead of requiring LLMs to directly predict precise physical parameters–a task challenging current LLMs’ capabilities–our method encodes physical properties into structured descriptors, explicitly guiding grasp pose prediction while avoiding error-prone low-level physics estimation.
@@ 3.3 Physical-Property-Oriented CoT Design
    $\texttt{obj}_{1}$>, $\texttt{obj}_{2}$>, …]. Total <#objs>.
@@
    The material properties of $\texttt{obj}_{\texttt{k}}$> are $\delta_{\texttt{phy}}$> (hardness), $\delta_{\texttt{phy}}$> (strength), and $\delta_{\texttt{phy}}$> (elasticity).
@@
    The appropriate verb to grasp the $\texttt{obj}_{\texttt{k}}$> is $\delta_{\texttt{act}}$>.
@@ 3.5 IntentGrasp Benchmark Construction
[…]
@@ 4.1 Main Results on IntentGrasp Benchmark
Setup. To evaluate the effectiveness of our grasp pose prediction method, we employ three metrics: coverage rate (CR) [26], earth mover’s distance (EMD) [41], and collision-free rate (CFR) [54] with appropriate adaptations. CR quantifies the proportion of ground-truth grasp poses that are sufficiently covered by predictions. Specifically, it measures the percentage of ground-truth poses where at least one predicted pose lies within a distance threshold $\theta$ (Euclidean distance in SE(3) space). We evaluate CR under three increasingly stringent thresholds: $\theta$ = 0.4, 0.3, and 0.2. EMD measures how holistically the predicted poses match the spatial distribution of ground-truth annotations in SE(3) space, where lower values indicate better distributional consistency. CFR evaluates the physical feasibility of predicted grasp poses by measuring their ability to avoid collisions with the target object and environment. However, a critical limitation arises when directly using the CFR metric: erroneous grasp poses far from the target object may artificially achieve higher scores because they avoid contact with environmental point clouds entirely. To jointly evaluates geometric consistency and physical feasibility where higher values indicate safer grasps, we propose an EMD-Weighted CFR (EW-CFR) metric defined as follows:
$\small\centering\text{EW-CFR}=\frac{1}{N}\sum_{i=1}^{N}\big{(}1-\mathcal{C}(\mathcal{P},\mathcal{M}(g_{i}))\big{)}\cdot\big{(}1-\frac{\text{EMD}(g_{i},\mathcal{\hat{G}})}{R}\big{)},\vskip-2.84544pt\@add@centering$ (3)
where $R$ normalizes the EMD($g_{i}$, $\mathcal{\hat{G}}$). The function $\mathcal{C}$ denotes a binary collision check between the point cloud $\mathcal{P}$ and the gripper model $\mathcal{M}(g_{i})$.
The IntentGrasp dataset is partitioned following conventional practice, with 80% of samples dedicated to training and the remaining 20% preserved as the evaluation set.
@@
Grasp Label Pruning. Wu [47] et al. demonstrate that excessively dense supervision incurs label ambiguity, particularly caused by the diversity of grasp orientations, as a critical bottleneck hindering model convergence. Our analysis reveals that certain objects in the Grasp-Anything-6D dataset contain over 700 annotated grasp poses. This overabundance of positive samples not only exacerbates the aforementioned issues but also severely challenges Hungarian matching during training. Consequently, we implemented a pruning strategy in IntentGrasp to reduce similar grasp labels, especially redundant rotation-similar annotations, per object to 100 instances. The detailed pruning method is provided in the supplementary material.
[…]
@@
Quantitative Results. In Tab. 1, our method presents superior performance across all metrics on the IntentGrasp benchmark under flexible language instructions. Notably, GraspCoT attains a CR of 0.5587 at the strictest threshold ($\theta$=0.2), outperforming 3DAPNet and LGrasp6D by a large margin of 0.2683 and 0.2238, respectively. This substantial improvement highlights our method’s ability to generate geometrically precise grasp poses even under high-precision requirements. The progressive performance gap across $\theta$=0.4, 0.3, and 0.2 further demonstrates that our approach maintains effectiveness as task difficulty escalates. The EMD of GraspCoT reflects a 14.1% relative reduction compared to LGrasp6D, indicating superior alignment between the predicted and ground-truth pose distributions. For our proposed EW-CFR metric, GraspCoT achieves a score of 0.4229, surpassing 3DAPNet and LGrasp6D by 20.1% and 9.9%, respectively. The w/o CoT version of GraspCoT confirms the critical role of our CoT reasoning pipeline: removing CoT results in a 25.4% decline in CR@0.2 and a 9.1% increase in EMD, respectively.
@@ 4.2 Ablation Studies
4.3 Real-World Experiment
(a) Real-world experimental setup.
(b) Example of the grasping task.
Figure 6: An application example on a Kinova Gen3 robotic arm. The video demonstration is included in supplementary materials.
Setup. We deploy a Kinova Gen3 robotic arm with an Intel RealSense D435i depth camera, as shown in Fig. 6. The system detects 6-DoF grasp poses from RGB-D data, maps them to end-effector poses via hand-eye calibration, and executes trajectories autonomously. Our ROS-based platform (NVIDIA RTX 3090) processes real-time camera data and controls the robot via a network interface. We categorize the test scenarios into explicit and flexible instructions, further dividing the latter into single-target grasping and multi-target sequential grasping. For each setting, we conduct 35 repeated grasping experiments, with the success rate as the evaluation metric. Partial successes (e.g., grasping one of two required targets) are counted as individual successes, with the rate defined as successful grasps / total targets.
Results. As shown in Tab. 5, when handling explicit grasping instructions, CoT-enhanced GraspCoT achieved a 54.2% success rate, outperforming its non-CoT counterpart by 2.8%. For flexible instructions, the CoT version maintained stable performance with a 5.6% advantage over the non-CoT variant, further validating its advantages in optimizing the grasp pose targeted at the object itself. Notably, in multi-target grasping scenarios where avoiding omissions of targets with weaker implicit relevance remains challenging, the CoT implementation improved success rates from 44.0% to 46.7%, which is a 2.7% enhancement, verifying its effectiveness in handling complex requirements.
@@ 5 Conclusion and Discussion
This work presents a novel 6-DoF grasp detection framework that bridges the gap between flexible instructions and robotic grasping. By integrating a physical-property-oriented CoT mechanism with QA-guided hierarchical reasoning, our method aligns target parsing, physical property analysis, and grasp action selection with grasp detection. The proposed multimodal LLM architecture, augmented by 3D-aware visual-textual token fusion, achieves state-of-the-art performance on a newly established benchmark IntentGrasp for diverse verbal-command grasping. Real-world validations further underscore its practical applicability.
While our method demonstrates robust performance under flexible instructions, some limitations persist. For instance, grasp predictions for slender objects (e.g., chopsticks) occasionally deviate from mass centers despite leveraging shape descriptors, suggesting the need to model latent physical properties beyond geometric cues. Additionally, future directions include extending the framework to sequential video inputs for dynamic scenes, such as grasping moving targets, which would require video-based grasping benchmarks to advance real-world adaptability.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62332016) and the Key Research Program of Frontier Sciences, CAS (No. ZDBS-LY-JSC001).